# First-passage properties of the jump process with a drift. The general case

### v1

@@ 4.2. Asymptotic behavior
Let us first consider the probability of the process staying positive up to the $n$-th jump. The Laplace transform of this probability
@@
In contrast to Sec. 4 it is more convenient to use explicit regularization for $\phi^{+}_{\text{a.c.}}(0;\rho,s)$ in (53), rather than utilizing (62) or treating the integral in (53) in the principal value sense. Specifically, we will use the following representation:
@@
The analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ in the $\lambda$-plane follows directly from those of $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$ as given in (55). In this section we focus on its behavior in the $s$-plane.
As will be shown shortly, the analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ depends on the sign of $\mu_{1}(0)=\langle M-\alpha t\rangle$, since $\mu_{1}(0)>0$ and $\mu_{1}(0)<0$ correspond to the survival and absorption regimes respectively, and $\mu_{1}(0)=0$ is a critical point. We will study these cases separately.
@@ 5.1. Survival regime
Figure 8. Schematic representation of the analytic structure of the integrand in (82) in the complex $k$-plane within the survival regime. The integrand features two branch points at $k=\mathrm{i}\zeta_{1,2}$ and two poles at $k=\mathrm{i}\epsilon$ and $k=-\mathrm{i}\lambda$. Upon continuation to $s>1$, the branch points pinch the shifted contour $\mathcal{C}_{*}$ inducing a singularity at $s=s_{*}$ defined by (85) in the $s$-plane.
Note that $c(0)=1$, and thus as $s\to 1$, we have $\zeta_{1}(0,s)\to 0$. This implies that the branch point at $\mathrm{i}\zeta_{1}(0,s)$ moves onto the contour of integration at $k=0$. Simultaneously, the pole at $k=\mathrm{i}\epsilon$ also approaches the contour at $k=0$ as $\epsilon\to 0$. This is exactly the coincidence of singularities studied in Sec. 4.
To perform the continuation to the region $s>1$, we shift the contour of integration in (82) from the real line to a horizontal contour $\mathcal{C}_{*}$ such that $\mathrm{Im}(k)=\zeta_{*}(0)$, where $\zeta_{*}(\rho)$ is the location of the minimum of $F(k;\rho)$ along the imaginary axis (see Fig. 4). Recall that if $\mu_{1}(0)>0$, then $\zeta_{*}(0)>0$, and when shifting the contour we must account for the residue at $k=\mathrm{i}\epsilon$. The resulting expression is
@@
Taking the limit $\epsilon\to 0$ is now straightforward and we have
@@
The analysis proceeds the same way. The only subtlety here is that the pole at $k=-\mathrm{i}\lambda$ lies in the upper half-plane. If $\lambda\in\left(-\zeta_{*},0\right)$, then while shifting the contour from the real line to $\mathcal{C}_{*}$ we must account for two poles: at $k=\mathrm{i}\epsilon$ and at $k=-\mathrm{i}\lambda$, with the contribution from the latter canceling out the prefactor.
@@
Summarizing, we showed that for $\lambda>-\zeta_{*}(0)$ the function $\hat{S}_{N}(s\,|\,\lambda)$ in the $s$-plane has a simple pole at $s=1$ and a branch point at $s=s_{*}$ with no other singularities for $\left|s\right|<s_{*}$. The schematic illustration is shown in Fig. 9. We emphasize that if $\lambda<-\zeta_{*}(0)$, then additional singularities may arise from the prefactor in (88). However, these singularities have no effect on the large-$n$ behavior of $S_{N}(n\,|\,X_{0})$, as will be explained shortly.
Figure 9. Schematic representation of the analytic structure of the $\hat{S}_{N}(n\,|\,\lambda)$ in the $s$-plane in the survival regime for $\lambda>-\zeta_{*}(0)$. There is a pole at $s=1$ and a branch point at $s_{*}$ defined by (86). In the unshaded region representations (82) and (88) are equivalent. In the shaded region they differ, with (88) providing a proper analytic continuation.
@@
The standard result in the context of analytic combinatorics [62] is that the large-$n$ behavior of the survival probability $S_{N}(n\,|\,X_{0})$ is governed by the singularities of its generating function closest to the origin. To find this generating function, we formally invert the Laplace transform with respect to $\lambda$ and find
@@
where the integration is performed along the vertical contour that lies to the right of all possible singularities of $\hat{S}_{N}(s\,|\,\lambda)$. For example, we can choose $\gamma$ such that $\gamma>-\zeta_{*}(0)$. With this choice, due to (88), the integrand $S_{N}(s\,|\,\lambda)$ has a pole at $s=1$ and a branch point at $s=s_{*}$. The pole corresponds to the asymptotic value of $S_{N}(n\,|\,X_{0})$ and the branching point at $s_{*}$ determines the rate at which this value is approached.
@@
Although the representation (92) is slightly different from (69) and (73) obtained in Sec. 4, their equivalence can be established by shifting the contour of integration in $\mathcal{G}_{N}(1\,|\,\lambda)$ as in (89) from $\mathcal{C}_{*}$ back to the real line. Since the detailed analysis of $\hat{S}_{\infty}(\lambda)$ and $S_{\infty}(X_{0})$ was already performed in Sec. 4, here we focus on the corrections for $S_{N}(n\,|\,X_{0})$.
@@
Recall that $s_{*}$ as in (86) does not depend on $\lambda$, hence the $n\to\infty$ behavior of the Laplace transform (93) translates into the $n\to\infty$ behavior of $S_{N}(n\,|\,X_{0})$, namely,
@@ 5.2. Absorption regime
We now proceed to the absorption regime in which $\mu_{1}(0)<0$ and the process eventually crosses the origin, $S_{\infty}(X_{0})=0$. The analysis is very similar to the one we have just performed in the survival regime. The main difference is that now $\zeta_{*}(0)<0$, and the contour of integration is shifted to the lower part of the $k$-plane. As a result, there is no contribution from the pole at $k=\mathrm{i}\epsilon$, hence, no pole at $s=1$ for $\hat{S}_{N}(s\,|\,\lambda)$, and the analytic continuation reads
$\alpha>\alpha_{c}:\quad\hat{S}_{N}(s\,|\,\lambda)=\left\{\begin{aligned} &\frac{1}{\lambda}\;\mathcal{G}_{N}(s\,|\,\lambda),\quad\lambda>-\zeta_{*}(0),\\ &\frac{1}{\lambda}\frac{\mathcal{G}_{N}(s\,|\,\lambda)}{1-sF\left(-\mathrm{i}\lambda;0\right)},\quad\lambda<-\zeta_{+}(0),\end{aligned}\right$. (96)
@@
Repeating the same arguments as for the survival regime we find that
@@ 5.3. Critical point
Finally, we analyze the behavior of $S_{N}(n\,|\,X_{0})$ at the critical point. Specifically we show that as the drift approaches the critical value the <<correlation length>> $\xi_{n}(\alpha)$ diverges resulting in the polynomial decay of the survival probability $S_{N}(n\,|\,X_{0})$ as $n\to\infty$. Additionally, we show that there is a nontrivial limiting behavior as $X_{0}\to\infty$ and $n\to\infty$ with fixed ratio $\frac{X_{0}}{\sqrt{n}}$ and compute the corresponding scaling function.
@@ 5.3.1. Divergence of the <<correlation length>>
where $\mathrm{i}\zeta_{*}\equiv\mathrm{i}\zeta_{*}(0)$ is the location of the minimum of $F(k;0)$ along the imaginary axis in the $k$-plane. If $\alpha=\alpha_{c}$, then this minimum is located exactly at the origin, $\zeta_{*}|_{\alpha=\alpha_{c}}=0$, which implies that $\frac{1}{\xi_{n}(\alpha_{c})}=0$; that is $\xi_{n}(\alpha)$ diverges as the strength of the drift approaches its critical value.
@@
When taking the average in the logarithm, the term with second derivative of $\zeta_{*}$ disappears and the expansion depends only on the first derivative. Specifically, we obtain
@@
By taking the derivatives with respect to $\alpha$ we can iteratively build the expansion of $\zeta_{*}(\alpha)$. Using that $\zeta_{*}|_{\alpha=\alpha_{c}}=0$ we obtain
@@
If $\alpha=\alpha_{c}$, then $s_{*}=1$ and $\xi_{n}(\alpha_{c})=\infty$, hence there is no exponential decay of $S_{N}(n\,|\,X_{0})$ as $n\to\infty$. On the level of the analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ in the $s$-plane this divergence is manifested by the coincidence of the pole and the branching point, as both of them are located at $s=1$. The large $n$ behavior of $S_{N}(n\,|\,X_{0})$ is governed by the behavior of $\hat{S}_{N}(s\,|\,\lambda)$ in the vicinity of $s=1$.
To construct this expansion we again resort to the representation (62) for $\phi^{\pm}(\lambda;\rho,s)$. Substituting (64) in (78) yields
@@
The analysis is similar to the one performed in Sec. 4 for the survival probability $S_{\infty}(X_{0})$. There is, however, a major difference. According to (74) the Laplace transform $\hat{S}_{\infty}(\lambda)$ has a singularity of the order $1/\lambda$, which yields a limiting behavior $S_{\infty}(X_{0})\to 1$ as $X_{0}\to\infty$. Extracting the subleading corrections requires a proper analytic continuation to $\mathrm{Re}(\lambda)<0$, which eventually results in (76). At the same time, as is clear from (109), when $\lambda\to 0$ the Laplace transform $\hat{U}(\lambda)$ of $U(X_{0})$ behaves as $1/\lambda^{2}$. Consequently, the leading order behavior for $X_{0}\to\infty$ can be extracted solely from the $\lambda\to 0$ expansion. While the continuation to $\mathrm{Re}(\lambda)<0$ can be straightforwardly performed utilizing the contour deformation in the $k$-plane, to simplify the discussion we do not present it here, especially since the resulting corrections are exponentially small as $X_{0}\to\infty$.
The asymptotic expansions of $U(X_{0})$ as $X_{0}\to 0$ and $X_{0}\to\infty$ are governed by the behaviors of $\hat{U}(\lambda)$ as $\lambda\to\infty$ and $\lambda\to 0$ respectively. The main technical challenge lies in expanding $I(\lambda;0,1)$. We give the detailed derivation in Appendix A, and here present the results. Specifically, substituting expansion (201) into (109) yields, for $X_{0}\to\infty$,
@@
$U(X_{0})\underset{X_{0}\to 0}{\sim}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha_{c}t)^{2}\rangle}\right]\right]\\ \times\left(1-\frac{X_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\log\left|1-F(k;0)\right|\right).$ (111)
@@
Before proceeding to the analysis of the survival probability $S_{T}(\tau\,|\,X_{0})$ two remarks are in order. First, the asymptotic expansions (110) and (111) are derived at the critical point $\alpha=\alpha_{c}$, under the assumption that the limit $n\to\infty$ is taken first. In this case, the survival probability $S_{N}(n\,|\,X_{0})$ decays algebraically as $1/\sqrt{n}$, with a prefactor depending on the initial position $X_{0}$. A natural question is, then, what happens when both $n$ and $X_{0}$ tend to infinity simultaneously. This leads to a non-trivial scaling limit, which is essentially equivalent to the Brownian motion.
Second, there exists a special case in which the jump distribution $q(M)$ and the waiting-time distribution $p(t)$ are tuned so that the effective random walk becomes symmetric. Although this situation is restrictive, it provides a valuable benchmark, since the survival probability of symmetric random walks on the half-line has been studied extensively. In this case, our expressions (108), (109), (110) and (111) should recover the results of [51].
@@
At the critical point $\alpha=\alpha_{c}$, the effective random walk consists of independent identically distributed jumps with zero mean and finite variance. In this situation it is natural to expect convergence to Brownian motion in the limit where both $X_{0}$ and $n$ tend to infinity with a fixed ratio $z=X_{0}/\sqrt{n}$. Accordingly, we anticipate that
@@
Next, we introduce the variable y=n(1s)y=n(1-s). In the limit $s\to 1$, the sum over $n$ can be replaced by an integral over $y$, while the survival probability $S_{N}(n\,|\,z\sqrt{n})$ can be replaced by its scaling form $V(z)$,
@@
This is precisely the equation studied in detail in [51] in the context of survival probabilities of one-dimensional random walks. Its solution reads
@@
It is worth emphasizing that the scaling behavior (118) does not rely on the specific structure of underlying effective random walk beyond the assumptions that it has zero mean, finite moments, and smooth density. In particular, the same functional form of the scaling limit was previously obtained in [51] for any symmetric random walk with finite variance. This essentially reflects the convergence of a random walk to a Brownian motion, which is rather a robust property, insensitive to microscopic details.
@@
Finally, let us consider the case where the effective random walk not only has a zero mean but also is symmetric. Since it requires fine-tuning of the distributions of the jump amplitudes $q(M)$ and inter-jump intervals $p(t)$, such a case is indeed very special. At the same time, the problem of survival probability of symmetric random walks was studied in great detail in [51]. In particular, the behavior of the survival probability of the symmetric random walk with light-tailed increment distribution was shown to be (in our notation) given by
@@
At first glance the representation (120) may seem to differ from (109), but they are actually the same. Note that the imaginary part of the integrand in (65) is odd, while the real part is even, hence by the parity argument, we can keep only the real part. Since for a symmetric random walk F(k;0)=F(k;0)F(k;0)=F(-k;0) we obtain

### v2

@@ 5.3. Numerical verification
Let us first consider the probability of the process staying positive up to the $n$-th jump. The Laplace transform of this probability, defined as
@@
In contrast to Section 5, it is more convenient to use explicit regularization for $\phi^{+}_{\text{a.c.}}(0;\rho,s)$ in (57), rather than utilizing (66) or treating the integral in (57) in the principal value sense. Specifically, we will use the following representation:
@@
The analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ in the $\lambda$-plane follows directly from that of $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$ as given in (59). In this section we focus on its behavior in the $s$-plane. We show that the Laplace transform (89) can be analytically continued following the same procedure as in the toy example of Section 3. In short, by shifting the integration contour in the $k$-plane from the real line to a horizontal line at $\mathrm{Im}(k)=\zeta_{*}$, we can perform the continuation up to the point $s_{*}$, where the two branch points of the logarithm coincide. This pinch singularity determines the decay rate $\xi_{n}$ in (14).
As will be shown shortly, the analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ depends on the sign of $\mu_{1}(0)=\langle M-\alpha t\rangle$, since $\mu_{1}(0)>0$ and $\mu_{1}(0)<0$ correspond to the survival and absorption regimes, respectively, and $\mu_{1}(0)=0$ is a critical point. Below, we study these cases separately. The resulting asymptotic expressions are then again tested against the numerical simulations with both inter-arrival times and jump amplitudes drawn from inverse-Gaussian distributions.
@@ 6.1. Survival regime
Figure 10. Schematic representation of the analytic structure of the integrand in (91) in the complex $k$-plane within the survival regime. The integrand features two branch points at $k=\mathrm{i}\zeta_{1,2}$ and two poles at $k=\mathrm{i}\epsilon$ and $k=-\mathrm{i}\lambda$. Upon continuation to $s>1$, the branch points pinch the shifted contour $\mathcal{C}_{*}=\{k:\,\mathrm{Im}(k)=\zeta_{*}(0)\}$ inducing a singularity at $s=s_{*}$ defined by (94) in the $s$-plane.
Note that $c(0)=1$, and thus as $s\to 1$, we have $\zeta_{1}(0,s)\to 0$. This implies that the branch point at $\mathrm{i}\zeta_{1}(0,s)$ moves onto the contour of integration at $k=0$. Simultaneously, the pole at $k=\mathrm{i}\epsilon$ also approaches the contour at $k=0$ as $\epsilon\to 0$. This is exactly the coincidence of singularities studied in Section 5.
To perform the continuation to the region $s>1$, we shift the contour of integration in (91) from the real line to a horizontal contour $\mathcal{C}_{*}$ such that $\mathrm{Im}(k)=\zeta_{*}(0)$ (see Fig. 10), where $\zeta_{*}(\rho)$ is the location of the minimum of $F(k;\rho)$ along the imaginary axis (see Fig. 6). Recall that if $\mu_{1}(0)>0$, then $\zeta_{*}(0)>0$, and when shifting the contour we must account for the residue at $k=\mathrm{i}\epsilon$. The resulting expression is
@@
Taking the limit $\epsilon\to 0$ is now straightforward, and we have
@@
The analysis proceeds in the same way. The only subtlety here is that the pole at $k=-\mathrm{i}\lambda$ lies in the upper half-plane. If $\lambda\in\left(-\zeta_{*},0\right)$, then while shifting the contour from the real line to $\mathcal{C}_{*}$ we must account for two poles: at $k=\mathrm{i}\epsilon$ and at $k=-\mathrm{i}\lambda$, with the contribution from the latter canceling out the prefactor.
@@
To summarize, we have shown that for $\lambda>-\zeta_{*}(0)$ the function $\hat{S}_{N}(s\,|\,\lambda)$ in the $s$-plane has a simple pole at $s=1$ and a branch point at $s=s_{*}$ with no other singularities for $\left|s\right|<s_{*}$. The schematic illustration is shown in Fig. 11. We emphasize that if $\lambda<-\zeta_{*}(0)$, then additional singularities may arise from the prefactor in (97). However, these singularities have no effect on the large-$n$ behavior of $S_{N}(n\,|\,X_{0})$, as will be explained shortly.
Figure 11. Schematic representation of the analytic structure of the $\hat{S}_{N}(s\,|\,\lambda)$ in the $s$-plane in the survival regime for $\lambda>-\zeta_{*}(0)$. There is a pole at $s=1$ and a branch point at $s_{*}$ defined by (95). In the unshaded region representations (91) and (97) are equivalent. In the shaded region they differ, with (97) providing a proper analytic continuation.
@@
The standard result in the context of analytic combinatorics [undefaam] is that the large-$n$ behavior of the survival probability $S_{N}(n\,|\,X_{0})$ is governed by the singularities of its generating function closest to the origin. To find this generating function, we formally invert the Laplace transform with respect to $\lambda$ and obtain
@@
where the integration is performed along the vertical contour that lies to the right of all possible singularities of $\hat{S}_{N}(s\,|\,\lambda)$. For example, we can choose $\gamma$ such that $\gamma>-\zeta_{*}(0)$. With this choice, due to (97), the integrand $S_{N}(s\,|\,\lambda)$ has a pole at $s=1$ and a branch point at $s=s_{*}$. The pole corresponds to the asymptotic value of $S_{N}(n\,|\,X_{0})$ and the branch point at $s_{*}$ determines the rate at which this value is approached.
@@
Although the representation (101) is slightly different from (73) and (77) obtained in Section 5, their equivalence can be established by shifting the contour of integration in $\mathcal{G}_{N}(1\,|\,\lambda)$ as in (98) from $\mathcal{C}_{*}$ back to the real line. Since the detailed analysis of $\hat{S}_{\infty}(\lambda)$ and $S_{\infty}(X_{0})$ was already performed in Section 5, here we focus on the corrections for $S_{N}(n\,|\,X_{0})$.
@@
Recall that $s_{*}$, as defined in (95), does not depend on $\lambda$, hence the $n\to\infty$ behavior of the Laplace transform (102) translates into the $n\to\infty$ behavior of $S_{N}(n\,|\,X_{0})$, namely,
@@ 6.2. Absorption regime
We now proceed to the absorption regime in which $\mu_{1}(0)<0$ and the process eventually crosses the origin, $S_{\infty}(X_{0})=0$. The analysis is very similar to that of the survival regime. The main difference is that now $\zeta_{*}(0)<0$, and the contour of integration is shifted to the lower part of the $k$-plane. As a result, there is no contribution from the pole at $k=\mathrm{i}\epsilon$, hence, there is no pole at $s=1$ for $\hat{S}_{N}(s\,|\,\lambda)$, and the analytic continuation reads
$\alpha>\alpha_{c}:\quad\hat{S}_{N}(s\,|\,\lambda)=\left\{\begin{aligned} &\frac{1}{\lambda}\;\mathcal{G}_{N}(s\,|\,\lambda),\quad\lambda>-\zeta_{*}(0),\\ &\frac{1}{\lambda}\frac{\mathcal{G}_{N}(s\,|\,\lambda)}{1-sF\left(-\mathrm{i}\lambda;0\right)},\quad\lambda\in(-\zeta_{+}(0),-\zeta_{*}(0)).\end{aligned}\right$. (105)
@@
Following the same procedure as in the survival regime, we find the result stated in (24), namely
@@ 6.3. Critical point
Finally, we analyze the behavior of $S_{N}(n\,|\,X_{0})$ at the critical point. Specifically, we show that as the drift approaches the critical value the <<correlation length>> $\xi_{n}(\alpha)$ diverges, resulting in polynomial decay of the survival probability $S_{N}(n\,|\,X_{0})$ as $n\to\infty$. Additionally, we show that there is a nontrivial limiting behavior as $X_{0}\to\infty$ and $n\to\infty$ with fixed ratio $\frac{X_{0}}{\sqrt{n}}$ and compute the corresponding scaling function.
@@ 6.3.1. Divergence of the <<correlation length>>
where $\mathrm{i}\zeta_{*}\equiv\mathrm{i}\zeta_{*}(0)$ is the location of the minimum of $F(k;0)$ along the imaginary axis in the $k$-plane. If $\alpha=\alpha_{c}$, then this minimum is located exactly at the origin, $\zeta_{*}|_{\alpha=\alpha_{c}}=0$, which implies that $\frac{1}{\xi_{n}(\alpha_{c})}=0$, that is, $\xi_{n}(\alpha)$ diverges as the strength of the drift approaches its critical value.
@@
When taking the average in the logarithm, the term with the second derivative of $\zeta_{*}$ disappears and the expansion depends only on the first derivative. Specifically, we obtain
@@
By taking the derivatives with respect to $\alpha$, we can iteratively build the expansion of $\zeta_{*}(\alpha)$. Using that $\zeta_{*}|_{\alpha=\alpha_{c}}=0$ we obtain
@@
If $\alpha=\alpha_{c}$, then $s_{*}=1$ and $\xi_{n}(\alpha_{c})=\infty$, hence there is no exponential decay of $S_{N}(n\,|\,X_{0})$ as $n\to\infty$. At the level of the analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ in the $s$-plane this divergence is manifested by the coincidence of the pole and the branch point, as both of them are located at $s=1$. The large-$n$ behavior of $S_{N}(n\,|\,X_{0})$ is governed by the behavior of $\hat{S}_{N}(s\,|\,\lambda)$ in the vicinity of $s=1$.
To construct this expansion, we again resort to the representation (66) for $\phi^{\pm}(\lambda;\rho,s)$. Substituting (68) in (87) yields
@@
The analysis is similar to the one performed in Section 5 for the survival probability $S_{\infty}(X_{0})$. There is, however, a major difference. According to (78) the Laplace transform $\hat{S}_{\infty}(\lambda)$ has a singularity of the order $1/\lambda$, which yields a limiting behavior $S_{\infty}(X_{0})\to 1$ as $X_{0}\to\infty$. Extracting the subleading corrections requires a proper analytic continuation to $\mathrm{Re}(\lambda)<0$, which eventually results in (80). At the same time, the Laplace transform $\hat{U}(\lambda)$ of $U(X_{0})$, as is clear from (118), behaves as $1/\lambda^{2}$ when $\lambda\to 0$. Consequently, the leading order behavior for $X_{0}\to\infty$ can be extracted solely from the $\lambda\to 0$ expansion. While the continuation to $\mathrm{Re}(\lambda)<0$ can be straightforwardly performed utilizing the contour deformation in the $k$-plane, to simplify the discussion we do not present it here, especially since the resulting corrections are exponentially small as $X_{0}\to\infty$.
The asymptotic expansions of $U(X_{0})$ as $X_{0}\to 0$ and $X_{0}\to\infty$ are governed by the behaviors of $\hat{U}(\lambda)$ as $\lambda\to\infty$ and $\lambda\to 0$, respectively. The main technical challenge lies in expanding $I(\lambda;0,1)$. We provide the detailed derivation in Appendix A, and present only the results here. Specifically, substituting expansion (223) into (118) yields, for $X_{0}\to\infty$,
@@
$U(X_{0})\underset{X_{0}\to 0}{\sim}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha_{c}t)^{2}\rangle}\right]\right]\\ \times\left(1-\frac{X_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\big[\log\left[1-F(k;0)\right]\big]\right).$ (120)
@@
Before proceeding to the analysis of the survival probability $S_{T}(\tau\,|\,X_{0})$, two remarks are in order. First, the asymptotic expansions (119) and (120) are derived at the critical point $\alpha=\alpha_{c}$, under the assumption that the limit $n\to\infty$ is taken first. In this case, the survival probability $S_{N}(n\,|\,X_{0})$ decays algebraically as $1/\sqrt{n}$, with a prefactor depending on the initial position $X_{0}$. A natural question is, then, what happens when both $n$ and $X_{0}$ tend to infinity simultaneously. This leads to a non-trivial scaling limit, which is essentially equivalent to the Brownian motion.
Second, there exists a special case in which the jump distribution $q(M)$ and the waiting-time distribution $p(t)$ are tuned so that the effective random walk becomes symmetric. Although this situation is restrictive, it provides a valuable benchmark, since the survival probability of symmetric random walks on the half-line has been studied extensively. In this case, our expressions (117), (118), (119) and (120) should recover the results of [undefaaa].
@@
At the critical point $\alpha=\alpha_{c}$, the effective random walk consists of independent identically distributed jumps with zero mean and finite variance. In this situation it is natural to expect convergence to the Brownian motion in the limit where both $X_{0}$ and $n$ tend to infinity with a fixed ratio $z=X_{0}/\sqrt{n}$. Accordingly, we anticipate that
@@
Next, we introduce the variable y=n(1s)y=n(1-s). In the limit $s\to 1$, the sum over $n$ can be replaced by an integral over $y$, while the survival probability $S_{N}(n\,|\,z\sqrt{n})$ can be replaced by its scaling form $V(z)$
@@
This is precisely the equation studied in detail in [undefaaa] in the context of survival probabilities of one-dimensional random walks. Its solution reads
@@
It is worth emphasizing that the scaling behavior (127) does not rely on the specific structure of the underlying effective random walk beyond the assumptions that it has zero mean, finite moments, and smooth density. In particular, the same functional form of the scaling limit was previously obtained in [undefaaa] for any symmetric random walk with finite variance. This essentially reflects the convergence of a random walk to a Brownian motion, which is a robust property, insensitive to microscopic details.
@@
Finally, let us consider the case where the effective random walk not only has a zero mean but is also symmetric. Since it requires fine-tuning of the distributions of the jump amplitudes $q(M)$ and inter-jump intervals $p(t)$, such a case is indeed very special. At the same time, the survival probability of symmetric random walks was studied in great detail in [undefaaa]. In particular, the behavior of the survival probability of the symmetric random walk with light-tailed increment distribution was shown, in our notation, to be given by
@@
At first glance, the representation (129) may seem to differ from (118), but they are actually the same. Note that the imaginary part of the integrand in (69) is odd, while the real part is even, hence, by the parity argument, we can keep only the real part. Since for a symmetric random walk F(k;0)=F(k;0)F(k;0)=F(-k;0) we obtain